We measured the electronics noise of the demodulation board, whitening board, and ADC for WFSs, and OPLEV board and ADC for DC QPD in MC2 transmission. We were using SR785.
Regarding the demodulation board, we did 2 series of measurements. For the first series of measurements, we were blocking WFS (attachment 1) and measuring noise at the output of the demod board (attachment 2a). This measurement includes dark noise of the WFS, electronics noise of demod board, and phase noise from LO. For the second series of the measurements, we were unplugging input to the demod board (attachment 2b & 2c is how they looked like before unplugging) (the mistake we made here is not putting 50-ohm terminator) and again measuring at the output of the demod board. This measurement doesn't include the dark noise of the WFS. We were measuring it for all 8 segments (I1, I2, I3, I4, Q1, Q2, Q3, Q4). The dark noise contribution is negligible with respect to demod board noise. In attachments 3 & 4 please find plots that include detection and demodulation contributions for both WFSs.
For whitening board electronics noise measurement, we were terminating the inputs (attachment 5) and measuring the outputs (attachment 6). Electronics noise of the whitening board is in the attachments 7 & 8.
For ADC electronics noise we terminated ADC input and measured noise using diaggui (attachments 9 & 10). Please find these spectra for WFS1, WFS2, and MC TRANS in attachments 11, 12 & 13.
For MC2 TRANS we measured OPLEV board noise. We did two sets of measurements, as for demod board of WFSs (with and without QPD dark noise) (attachments 14, 15 & 16). In the case of OPLEV board noise without dark noise, we were terminating the OPLEV input. Please find the electronics noise of OPLEV's segment 1 (including dark noise which is again much smaller with respect to the OPLEV's electronics noise) in attachment 17.
For the transfer functions, demod board has flat tf, whitening board tf please find in attachment 18, ADC tf is flat and it is (2**16 - 1)/20 [cts/V], and dewhitening tf please find in attachment 19. Also please find the ASD of the spectral analyzer noise (attachment_20).
Measurements for WFS1 demod and whitening were done on 5th of July between 15h and 18h local time. Measurements for WFS2 demod and whitening were done on 6th of July between 15h and 17h local time. All the rest were done on July 7th between 14h and 19h. In attachment 21 also find the comparison between electronics noise for WFSs and cds error signal (taken on the 28th of June between 17h and 18h). Sorry for bad quality of some pictures.